Using High-Low to Calculate Fixed Cost, Calculate the Variable Rate, and Construct a Cost Function Pizza Vesuvio makes specialty pizzas. Data for the past 8 months were collected:

Month Labor Cost($) Employee Hours

January 9,290 510

February 6,800 590

March 7,331 620

April 7,940 470

May 9,587 550

June 8,290 430

July 12,508 700

August 7,300 390

Pizza Vesuvio's controller wants to calculate the fixed and variable costs associated with labor used in the restaurant. In your calculations, round the variable rate per employee hour to the nearest cent. If required, round your final answers to the nearest cent. Required:

1. Using the high-low method, calculate the fixed cost of labor. $ 2,141

2. Using the high-low method, calculate the variable rate. $ 18.41 per employee hour

3. Using the high-low method, construct the cost formula for total labor cost. Total labor cost = $ 12,508 + ($ 18.41 × Employee hours)
Answer #1

1) Fixed cost = 12508-(700*16.80) = 748

2) Variable rate = Change in Cost/Change in Hour = (12508-7300)/(700-390) = 16.80 per employee hour

3) Formula :

Y = a+bx

Y = 748+16.80X